OpenStudy (anonymous):

simplify the expression if not possible write simplified but dont confuse me lol :P 2(11d-5)

OpenStudy (anonymous):

22d-10* 11X2=22 -5X2=-10

OpenStudy (anonymous):

u confused meh

OpenStudy (anonymous):

what part?